f(x)=x+1/xsquared+x -6 find x and y intercepts and the domain
Answer by stanbon(75887)

f(x)=x+1/xsquared+x -6 find x and y intercepts and the domain
-----
f(x) = (x+1)/(x^2+x-6)
-----
Factor:
f(x) = (x+1)/[(x+3)(x-2)]
---
x-intercept: let y = 0 and solve for "x":
y = 0 when the numerator of the fraction is zero,
i.e. when x= -1.
------------
y-intercept: Let x = 0 and solve for "y":
y = 1/(-6) = -1/6
--------------------
Domain: The denominator cannot be zero:
x cannot be -3 or 2
Domain is All Real Numbers except x = -3 or 2.
=================
Cheers,
Stan H.